OpenStudy (anonymous):

Determine the number of solutions of 2x2 + 3x = 0. A. 0 B. 1 (double root) C. 2 i got A

OpenStudy (anonymous):

@e.cociuba

OpenStudy (e.cociuba):

How did u get 0?? b^2 - 4ac = 9 - 4 * 2 * 0 = 9 > 0 two roots Answer: c. 2 :)